OpenStudy (anonymous):

An animal shelter is deciding between two brands of cat food. Two tubs of cat food are shown. The first one is Happy Cat. It costs forty-eight dollars for fifteen pounds. The second one is Cat's Meow. It costs fifty dollars for sixteen pounds. The shelter is happy with both brands of cat food, so they will make their decision based on price. Which brand of cat food is the better buy? Explain how you know, showing all your work.

OpenStudy (anonymous):

Just divide to find unit price: Happy Cat: $48/ 15 lb = ... Cat Meou : $50/ 16lb = ... Which one is cheaper?

OpenStudy (anonymous):

48/15=3.2 50/16 = 3.125

OpenStudy (anonymous):

Now decide it yourself which one is better buy?

OpenStudy (anonymous):

3.2